# A Lesson in Splats: Teacher-Guided Diffusion for 3D Gaussian Splats Generation with 2D Supervision

### v2

@@ Abstract
We introduce a diffusion model for Gaussian Splats, SplatDiffusion, to enable generation of three-dimensional structures from single images, addressing the ill-posed nature of lifting 2D inputs to 3D. Existing methods rely on deterministic, feed-forward predictions, which limit their ability to handle the inherent ambiguity of 3D inference from 2D data. Diffusion models have recently shown promise as powerful generative models for 3D data, including Gaussian splats; however, standard diffusion frameworks typically require the target signal and denoised signal to be in the same modality, which is challenging given the scarcity of 3D data. To overcome this, we propose a novel training strategy that decouples the denoised modality from the supervision modality. By using a deterministic model as a noisy teacher to create the noised signal and transitioning from single-step to multi-step denoising supervised by an image rendering loss, our approach significantly enhances performance compared to the deterministic teacher. Additionally, our method is flexible, as it can learn from various 3D Gaussian Splat (3DGS) teachers with minimal adaptation; we demonstrate this by surpassing the performance of two different deterministic models as teachers, highlighting the potential generalizability of our framework. Our approach further incorporates a guidance mechanism to aggregate information from multiple views, enhancing reconstruction quality when more than one view is available. Experimental results on object-level and scene-level datasets demonstrate the effectiveness of our framework.
Figure 1: (Left) Standard diffusion training is typically constrained to same-modality supervision. We introduce a framework that breaks this barrier by decoupling the sources of noised samples and supervision. Leveraging imperfect predictions of a pretrained feedforward 3D reconstruction module, we train a denoiser that operates in 3D, while being supervised by sparse 2D views. (Right) When paired with two different noisy teachers, our diffusion model enhances reconstruction quality across both objects and scenes, particularly in regions where deterministic models struggle due to reconstruction ambiguity.
@@ 1 Introduction
3D reconstruction from single 2D images is essential for computer vision applications, such as augmented reality, robotics, and autonomous vehicles, which rely on inferring 3D structures from limited viewpoints. Lifting 2D images to 3D is an ill-posed problem, because different 3D shapes can produce identical 2D projections.
Current approaches for 3D reconstruction from single images can be categorized into two main types: deterministic predictions and generative models, each with distinct limitations that stem from the ill-posed nature of inferring 3D structures from 2D images.
@@
A prevalent approach in 3D reconstruction is to use deterministic neural networks to map input images to 3D representations, such as Neural Radiance Fields (NeRF) [33], 3D Gaussian Splats (3DGS) [39], and triplanes [68]. By leveraging differentiable rendering techniques, these methods can train directly from 2D images, circumventing the need for large volumes of 3D data, which are relatively scarce. This is advantageous because annotated 3D data is often difficult or impractical to obtain, especially for real-world applications. However, despite ongoing performance improvements, deterministic models remain inherently limited by the ambiguity in the 2D-to-3D mapping. These models cannot fully capture the range of possible 3D structures that correspond to a single 2D image, leading to overly smooth or blurred outputs when supervised by appearance-based losses for a specific target. This limitation constrains their ability to generate diverse and accurate 3D reconstructions, particularly when multiple plausible 3D structures could correspond to the same 2D input.
@@
In contrast, generative models, especially diffusion models [15, 8], have recently shown strong potential in representing 3D data. Diffusion models are trained to progressively denoise corrupted versions of 3D data to generate 3D outputs that are likely under the training set distribution, either by directly operating in the 3D space [1, 30, 34] or in a lower-dimensional latent space [40, 51, 39]. These models have proven powerful for generating realistic data across various modalities. However, diffusion models for 3D generation face a fundamental limitation due to their training process, in which the denoising model is trained on noisy samples using their clean counterparts as supervision. This requirement demands a substantial amount of 3D data, making these models difficult to scale to real-world applications where 3D data is limited. Consequently, while diffusion models are highly effective, their dependency on large 3D datasets restricts their applicability for 3D generation when starting from 2D image data.
@@
Some attempts have been made to bypass these limitations by training 3D generative models using multi-view images [46]. These models aggregate information across multiple views, structuring predictions in 3D space. However, they still fall short in generation quality compared to deterministic methods and require multiple views for each target object, which is often impractical in real-world applications. Thus, although both deterministic and generative models have made strides in 3D reconstruction, the field lacks scalable, high-performance solutions that can infer 3D structures from single 2D images. Research into training 3D diffusion models from single-image inputs remains underexplored, highlighting an important gap that our work aims to address.
In this work, we propose a novel training strategy that fundamentally revises the principles of diffusion model training by decoupling the denoised modality from the supervision modality. To clarify the challenge, we note that although a denoiser can be trained in the 3D modality and supervised with 2D images, diffusion models require the noisy signal and the clean signal to remain in the same modality—here, in 3D. However, since our supervision comes from 2D images, the key question becomes: how to create a noisy 3D signal suitable for denoising?
Our solution leverages powerful deterministic 3D reconstruction methods as “noisy teachers”. While deterministic models can produce 3D representations that are imperfect, as manifested by blurry renderings, they nonetheless provide a useful starting point in 3D. This approach is advantageous because the goal in diffusion training is to apply noise to the signal. By introducing enough noise beyond a critical timestep $t^{*}$, the difference between the noisy 3D signal provided by the deterministic model and the true, clean 3D structure becomes minimal. This "sweet spot" in noise level, inspired by techniques like SDEdit [60], allows us to use the deterministic 3D output to create a suitable noisy input for diffusion training. At this stage, we can train a denoiser to predict the clean signal as is often done for guidance using Tweedie’s formula [10, 22, 38], and we can supervise this prediction using image-based rendering losses.
@@
However, this alone is not sufficient because if the denoiser only learns from timesteps $t>t^{*}$, it is bound to producing blurry outputs. To overcome this, we introduce a second key innovation: a multi-step denoising strategy that replaces the traditional single-step denoising framework. Specifically, starting from a noise level $t>t^{*}$, our model performs iterative denoising, akin to its behavior during inference, progressively reducing noise over multiple steps until reaching the clean 3D structure. This multi-stage approach allows the model to generate sharp and detailed reconstructions, as each denoising step moves closer to the true 3D target. By supervising the final output with image-based rendering losses, we enable the model to learn to produce high-quality outputs while updates propagate across the entire denoising sequence. Consequently, our approach opens up the capability to directly train 3D diffusion models from 2D supervision.
@@
Notably, this strategy is extremely flexible and can utilize various 3D teacher models as long as the diffusion prediction can be supervised by them. In our experiments, we demonstrate this flexibility using two types of deterministic models: the popular Splatter Image [48] and the more advanced and recent Flash3D [47]. With these models, we train on single object and scene data, respectively. In both cases, our method significantly improves the performance of the base teacher model by $0.5-0.85$ PSNR. Additionally, our diffusion model facilitates the incorporation of additional views through guidance, further boosting performance compared to standard optimization.
@@ 2 Related Work
Diffusion Models Trained Directly on 3D Data. One line of research focuses on designing diffusion models that directly operate in 3D space. These models have been developed for various 3D representations, including point clouds [51, 31, 65], meshes [1, 30], 3D Gaussian splats [34, 39], and neural fields [7, 35, 9, 43, 4]. While effective, these methods assume the availability of high-quality 3D datasets in the target representation, which are often scarce and lack the breadth of real-world diversity. This data scarcity limits the generalization and applicability of these models, particularly in in-the-wild scenarios.
@@
In contrast, our approach minimizes the distribution discrepancy between teacher-induced noised samples and (unavailable) target noise samples by focusing on large noise values and refining lower-noise predictions through a multi-step denoising process. Several approaches [2, 49, 46], denoise multi-view images using a denoiser structured to predict a 3D representation, which is then rendered into 2D views. However, these methods inherently rely on the bijectivity of multi-view and 3D representations, which only hold with a substantial number of images. Additionally, because the images are noised independently, they may not coherently represent the noisy 3D structure, potentially harming consistency. Our proposed method, in contrast, directly denoises within the 3D representation while using 2D views for supervision, addressing both data scarcity and view coherence by explicitly working in 3D space.
@@ 3 Method
Problem Formulation. We tackle the problem of training a 3D diffusion model, $D_{\theta}(s_{t},t,x_{\text{src}})$, that maps $N$ noisy 3D Gaussian Splats, $s_{t}\in\mathbb{R}^{N\times d}$ to their clean version $s_{0}$. Each of the Gaussian Splats is of dimension $d$, representing properties such as center, covariance, opacity, and color. The model is conditioned on a single image $x_{\text{src}}$ and uses $k\geq 1$ additional views of the same content for supervision, $\{x_{\text{tgt}}^{v}\}_{v=0}^{k-1}$, without access to 3D ground truth. We assume access to a pre-trained deterministic model $s_{0}^{\text{teacher}}=T_{\phi}(x)$, trained on the same sparse view data, that reconstructs 3D Gaussian Splats from a single image—or we can train such a model ourselves. Our method employs this trained model as a noisy teacher, generating noisy samples to train the diffusion model, which is supervised by the target image set $\{x_{\text{tgt}}^{v}\}_{v=0}^{k-1}$.
Overview. Our pipeline operates in two stages. First, we bootstrap the diffusion model by supervising it with the noisy teacher’s predictions (Section 3.2). We then proceed to fine-tune the diffusion model using multi-step denoising and rendering losses (Section 3.1). Both stages are further equipped with a cycle consistency regularization described in Section 3.3. Although the bootstrapping stage precedes fine-tuning in the pipeline, we present it second in this manuscript to facilitate a smoother explanation of our core contributions. The model pipeline is depicted in Fig. 2.
@@ 3.1 Decoupling Noised Samples from Supervision with Multi-Step Denoising
Our approach to overcoming the aforementioned uni-modality limitation of diffusion model training is to decouple the source for the noisy samples from the supervision. Specifically, in standard diffusion training, noise is added to the target ground truth sample, which is then fed to the denoiser for recovering the clean target. Here, we do not have access to true 3D target data; instead, we replace it with a 3D prediction from a pretrained deterministic model. As previously discussed this model is limited in its ability to generate the diverse plausible 3D structures often resulting in blurry and imprecise predictions, thus we consider it to be a “noisy teacher”. A key insight is that while the noisy teacher does not produce 3D Gaussian Splats (3DGS) that are sufficient as a standalone solution, they are useful as a starting point in our proposed framework. We further take inspiration from [60], which finds that with enough noise, the data distribution of two modalities can overlap. Based on this, we choose a timestep $t^{*}$ such that for $t\geq t^{*}$, the noisy samples generated by the noisy teacher are likely to align with those that would have resulted from a forward noising process applied to the true, unknown ground truth 3DGS. Denoting these samples as
@@ 3.2 Noisy Teacher Bootstrapping
Training a 3D diffusion model directly using the multi-step denoising paradigm described above is computationally expensive. This is primarily due to the increased memory costs of maintaining gradients over multiple denoising steps in 3D space, which limits batch sizes and reduces efficiency. To address this, we propose avoiding this training approach from scratch by first bootstrapping our model using the noisy teacher.
@@ 4 Experiments
Splatter Image [48] 24.00 0.92 0.078 24.43 0.93 0.067
SplatDiffusion 24.84 0.93 0.077 25.21 0.93 0.066
@@
4.1 Experimental Setups
Datasets. We conduct experiments using two datasets: the object-level ShapeNet-SRN [6, 44] and the scene-level RealEstate10k [66]. ShapeNet-SRN comprises synthetic objects across various categories. In line with Splatter Image [48] and PixelNeRF [61], we focus on the cars and chairs classes. The resolution for ShapeNet-SRN dataset is $128\times 128$, and the Splatter Image model is employed as the teacher for the ShapeNet experiments. RealEstate10k consists of real-world video data captured in both indoor and outdoor environments. Following Flash3D [47], we use a resolution of $256\times 384$ for training in our experiments. The Flash3D model serves as the teacher to guide our diffusion model at the bootstrapping stage.
Method Memory Usage (GB) Model Size (MB)
PixelNeRF [61] 3.05 113
VisionNeRF [27] 6.42 1390
Splatter Image [48] 1.71 646
Ours 1.15 295
Table 2: Memory Footprint and Model Size.
Memory Usage and Model Size We report both GPU memory consumption and model size. As shown in Tab.2, our model exhibits a significantly smaller size compared to VisionNeRF and Splatter Image. While PixelNeRF has a smaller model size, our approach achieves lower GPU memory consumption on the ShapeNet-SRN dataset.
4.2 Implementation Details.
Multi-step Denoising. We train the model using 4 NVIDIA A6000 GPUs. The computational efficiency is demonstrated in Tab. 2. During the bootstrapping stage (stage 1), a batch size of $100$ per GPU is employed to train the diffusion model under the guidance of the teacher model. Following this, in stage 2, multi-step denoising is performed using a DDIM sampler with 10 inference steps. To manage the increased computational complexity during this phase, the batch size is reduced to $10$.
Misc. We use Adam [23] as our optimizer with $\beta_{1}=0.9,\beta_{2}=0.999$. We use a total noising steps of $100$, with a linear scheduling, starting from $0.0001$ to $0.2$. For the bootstrapping stage, we use the teacher model to provide both supervision and noised samples. Instead of predicting the noises added to the splatters, our diffusion model denoises the noised inputs to clean samples directly. We set $t^{*}$ to be 20. For the architecture of diffusion model, we use the U-Net implementation from diffusers  [^1] https://huggingface.co/docs/diffusers. For the consistency branch, we use a denoising step of 10.

### v3

@@ Abstract
We present a novel framework for training 3D image-conditioned diffusion models using only 2D supervision. Recovering 3D structure from 2D images is inherently ill-posed due to the ambiguity of possible reconstructions, making generative models a natural choice. However, most existing 3D generative models rely on full 3D supervision, which is impractical due to the scarcity of large-scale 3D datasets. To address this, we propose leveraging sparse-view supervision as a scalable alternative. While recent reconstruction models use sparse-view supervision with differentiable rendering to lift 2D images to 3D, they are predominantly deterministic, failing to capture the diverse set of plausible solutions and producing blurry predictions in uncertain regions. A key challenge in training 3D diffusion models with 2D supervision is that the standard training paradigm requires both the denoising process and supervision to be in the same modality. We address this by decoupling the noisy samples being denoised from the supervision signal, allowing the former to remain in 3D while the latter is provided in 2D. Our approach leverages suboptimal predictions from a deterministic image-to-3D model—acting as a "teacher"—to generate noisy 3D inputs, enabling effective 3D diffusion training without requiring full 3D ground truth. We validate our framework on both object-level and scene-level datasets, using two different 3D Gaussian Splat (3DGS) teachers. Our results show that our approach consistently improves upon these deterministic teachers, demonstrating its effectiveness in scalable and high-fidelity 3D generative modeling. See our project page https://lesson-in-splats.github.io/.
Figure 1: (Left) Standard diffusion training is constrained to same-modality supervision. We break this barrier by decoupling the sources of noised samples and supervision. Leveraging imperfect predictions of a feedforward 3D reconstruction module, our method offers a fully image-based 3D diffusion training scheme. (Right) When paired with two different noisy teachers, our diffusion model enhances reconstruction quality and 3D geometry across both objects and scenes. Notably, our model is trained on the same data as the teachers, and uses a smaller model size. "Medium" and "Large" denote the model size, see Sec. 4.1.
@@ 1 Introduction
3D Reconstruction is essential for computer vision applications, such as augmented reality, robotics, and autonomous driving [30, 13, 8, 42], which rely on inferring 3D structures from limited viewpoints. However, reconstructing 3D objects or scenes from 2D images is challenging. First, it is an ill-posed problem because different 3D shapes can produce identical 2D projections. Second, 3D datasets are scarce, especially in comparison to their image dataset counterparts, limiting the ability to directly train on 3D data. Current approaches for 3D reconstruction from single images can be categorized into two main types: deterministic predictions and generative models, each with distinct limitations.
@@
A prevalent approach in 3D reconstruction is to use deterministic feedforward neural networks to map input images to 3D representations, such as Neural Radiance Fields (NeRF) [37, 19] and 3D Gaussian Splats (3DGS) [23, 55, 54, 70]. Leveraging differentiable rendering techniques, these methods can be trained directly from sprase 2D views, circumventing the need for large volumes of 3D data. This is advantageous because 3D data is often difficult or impractical to obtain, especially for real-world applications. However, despite ongoing performance improvements, deterministic models remain inherently limited by the ambiguity in the 2D-to-3D mapping. These models cannot fully capture the range of possible 3D structures that correspond to a source image, leading to overly smooth or blurred outputs when supervised by appearance-based losses.
@@
In contrast, diffusion models [17, 10], have recently shown a strong potential in generating 3D data. 3D Diffusion models are trained to progressively denoise corrupted versions of 3D data to generate 3D outputs that are likely under the training set distribution, either by directly operating in the 3D space [1, 33, 38, 34, 72] or in a higher-dimensional latent space  [47, 58, 46]. However, diffusion models for 3D generation face a fundamental limitation due to their training process, in which the denoiser – which operates in 3D – is trained on noisy samples using their clean counterparts as supervision. This requirement demands a substantial amount of 3D data, making these models difficult to scale to real-world applications where 3D data is limited. Some attempts have been made to bypass these limitations by training 3D generative models using multi-view images [53, 66]. These models aggregate information across multiple views, structuring predictions in 3D space. However, such methods rely on the bijectivity of multi-view and 3D representations, which only holds for a substantial number of images limiting their applicability. When the number of views is limited, they often fall short in generation quality.
@@
Thus, although both deterministic and generative models have made strides in 3D reconstruction, the field lacks scalable, high-performance solutions that can infer 3D structures from single 2D images. Research into training 3D diffusion models using only 2D supervision remains underexplored, highlighting an important gap that our work aims to address.
In this work, we propose a novel training strategy that fundamentally revises the principles of diffusion model training by decoupling the denoised modality (3D) from the supervision modality (2D). This stands in contrast to traditional diffusion training, which requires the noisy and clean signals to remain in the same modality—here, in 3D. Our solution leverages deterministic 3D reconstruction methods as “noisy teachers”. While deterministic 3D predictions are imperfect and exhibit artifacts, we show that they nonetheless can generate useful 3D samples as input to the denoiser. Specifically, by introducing noise beyond a critical timestep $t^{*}$, the noisy 3D signal provided by the deterministic model nearly matches that of the (unavailable) true 3D structure. This “sweet spot” in noise level is inspired by techniques like SDEdit [36]. Once denoised, the predicted clean 3D structure can be rendered and supervised with reference images, alleviating the need for 3D supervision.
@@
However, this alone is not sufficient because if the denoiser only learns from timesteps $t>t^{*}$, it is bound to produce blurry outputs, thus not able to fully exploit fine-grained details in images. To overcome this, we introduce a second key innovation: a multi-step denoising strategy that replaces the traditional single-step denoising framework. Specifically, starting from a noise level $t>t^{*}$, our model performs iterative denoising, akin to its behavior during inference, progressively reducing noise over multiple steps until reaching its sharpest 3D estimate at $t=0$. Rendering this output, enables supervising the model with ground truth images, effectively propagating gradients across lower time steps $t\leq t^{*}$ to adapt the denoiser to generate high quality reconstructions. In summary, by leveraging 2D supervised deterministic teachers, and multi-step denoising, our method offers a fully image-based 3D diffusion training scheme.
@@
Notably, this strategy is flexible and can utilize various teacher models. In our experiments, we demonstrate this flexibility using two types of deterministic models: Splatter Image [55] and Flash3D [54]. With these models, we train on single object and scene data, respectively. In both cases, our method significantly improves the performance of the base teacher model by 0.50.850.5-0.850.5 - 0.85 PSNR. Additionally, our diffusion model facilitates the incorporation of additional views through the guidance mechanism, further boosting performance compared to standard optimization.
@@ 2 Related Work
Diffusion Models Trained Directly on 3D Data. One line of research focuses on designing diffusion models that directly operate in 3D space. These models have been developed for various 3D representations, including point clouds [58, 34, 72, 40, 30, 29], meshes [1, 33], 3D Gaussian splats [38, 46, 43], and neural fields [7, 39, 11, 50, 4, 41, 9]. While effective, these methods assume the availability of high-quality 3D datasets in the target representation, which are often scarce and lack the breadth of real-world diversity. This data scarcity limits the generalization and applicability of these models, particularly in in-the-wild scenarios.
@@
In contrast, our approach minimizes the distribution discrepancy between teacher-induced noised samples and (unavailable) target noise samples by focusing on large noise values and refining lower-noise predictions through a multi-step denoising process. Several approaches [2, 56, 53, 66], denoise multi-view images using a denoiser structured to predict a 3D representation, which is then rendered into 2D views. However, these methods inherently rely on the bijectivity of multi-view and 3D representations, which only holds with a substantial number of images. Additionally, because the images are noised independently, they may not coherently represent the noisy 3D structure, potentially harming consistency. Our proposed method, in contrast, directly denoises within the 3D representation while using 2D views for supervision, addressing both data scarcity and view coherence by explicitly working in 3D space.
@@ 3 Method
Problem Formulation. We tackle the problem of training an image conditioned 3D diffusion model from 2D views only. A denoiser $D_{\theta}(s_{t},t,x_{\text{src}})$, maps $N$ noisy 3D Gaussian Splats, $s_{t}\in\mathbb{R}^{N\times d}$ to their clean version $s_{0}$. Each of the Gaussians is of dimension $d$, representing properties such as center, covariance, opacity, and color. The model is conditioned on a single image $x_{\text{src}}$ and uses $k\geq 1$ additional views of the same content for supervision, $\{x_{\text{tgt}}^{v}\}_{v=0}^{k-1}$, without access to 3D ground truth. We assume access to a pre-trained deterministic model $s_{0}^{\text{teacher}}=T_{\phi}(x)$, trained on the same sparse view data, that reconstructs 3D Gaussian Splats from a single image—or we can train such a model ourselves. Our method employs this trained model as a noisy teacher, generating noisy samples to train the diffusion model, which is supervised by the target image set $\{x_{\text{tgt}}^{v}\}_{v=0}^{k-1}$.
Overview. Our pipeline operates in two stages. First, we bootstrap the diffusion model by supervising it with the noisy teacher’s predictions (Section 3.2). We then proceed to fine-tune the diffusion model using multi-step denoising with rendering losses (Section 3.1). Both stages are further equipped with a cycle consistency regularization described in Section 3.3. Although the bootstrapping stage precedes fine-tuning in the pipeline, we present it second in this manuscript to facilitate a smoother explanation of our core contributions. The model pipeline is depicted in Fig. 2.
@@ 3.1 Decoupling Noised Samples from Supervision with Multi-Step Denoising
Our approach to overcoming the aforementioned uni-modality limitation of diffusion model training is to decouple the source for the noisy samples from the supervision. Specifically, in standard diffusion training, noise is added to the target ground truth sample, which is then fed to the denoiser for recovering the clean target. Here, we do not have access to true 3D target data; instead, we replace it with a 3D prediction from a pretrained deterministic model. As previously discussed this model is limited in its ability to generate the diverse plausible 3D structures often resulting in blurry and imprecise predictions, thus we consider it to be a “noisy teacher”. A key insight is that while the noisy teacher does not produce 3D Gaussian Splats (3DGS) that are sufficient as a standalone solution, they are useful as a starting point in our proposed framework. We further take inspiration from SDEdit [36], which finds that with enough noise, the data distribution of two modalities can overlap. Based on this, we choose a timestep $t^{*}$ such that for $t\geq t^{*}$, the noisy samples generated by the noisy teacher are likely to align with those that would have resulted from a forward noising process applied to the true, unknown ground truth 3DGS. Denoting these samples as
@@ 3.2 Noisy Teacher Bootstrapping
Training a 3D diffusion model directly using the multi-step denoising paradigm is computationally expensive. This is due to the increased memory costs of maintaining gradients over multiple denoising steps in 3D space, which limits batch sizes and reduces efficiency. To address this, we propose avoiding this training approach from scratch by first bootstrapping our model using the noisy teacher.
@@ 4 Experiments
Splatter Image (Large) [55] 24.00 0.92 0.078 24.43 0.93 0.067
SplatDiffusion (Medium) 24.84 0.93 0.077 25.21 0.93 0.066
@@
4.1 Experimental Setups
Memory Usage and Model Size. Due to limited computational resources, our diffusion model utilizes a smaller U-Net architecture (Medium) compared to the original Splatter Image model (Large). In our ablation studies, we train a Splatter Image using our "Medium" U-Net and report its performance. Unless stated otherwise, all experiments report the performance of the original Splatter Image model (Large), which serves as a teacher for our smaller model (Medium).
We report both GPU memory consumption and model size in Tab.3. Our model exhibits a significantly smaller size compared to VisionNeRF and Splatter Image. While PixelNeRF has a smaller model size, our approach achieves lower GPU memory consumption on the ShapeNet-SRN dataset.
Datasets. We conduct experiments using two datasets: the object-level ShapeNet-SRN [6, 51] and the scene-level RealEstate10k [73]. ShapeNet-SRN comprises synthetic objects across various categories. In line with Splatter Image [55] and PixelNeRF [68], we focus on the cars and chairs classes. The resolution for ShapeNet-SRN dataset is $\times 128$ 128, and the Splatter Image model is employed as the teacher for the ShapeNet experiments. RealEstate10k consists of real-world video data captured in both indoor and outdoor environments. Following Flash3D [54], we use a resolution of 256×384256$\times$ 384256 × 384 for training in our experiments. The Flash3D model serves as the teacher to guide our diffusion model at the bootstrapping stage.
Method Memory Usage (GB) Model Size (MB)
PixelNeRF [68] 3.05 113
VisionNeRF [28] 6.42 1390
Splatter Image (Large) [55] 1.71 646
Ours (Medium) 1.15 295
Table 3: Memory Footprint and Model Size.
4.2 Implementation Details
Multi-step Denoising. We train the model using 4 NVIDIA A6000 GPUs. The computational efficiency is demonstrated in Tab. 3. During the bootstrapping stage (stage 1), a batch size of 100100100 per GPU is employed to train the diffusion model under the guidance of the teacher model. Following this, in stage 2, multi-step denoising is performed using a DDIM sampler with 10 inference steps. To manage the increased computational complexity during this phase, the batch size is reduced to 101010.
Further implementation details are provided in the appendix.